OpenStudy (anonymous):

solve for x: 9(10^x-2)=23?

OpenStudy (anonymous):

we have to use log base 10 to get x. 10^x - 2 = 23/9 10^x = 2 + 23/9 Take log both side \[\log_{10}10^{x} = \log_{10}\left( 2 + 23/9 \right) \]\[x*\log_{10}10 = \log_{10}\left( 2 + 23/9 \right) \]We know that \[\log_{10}10 = 1 \] hence \[x=\log_{10}\left( 2 + 23/9 \right) \] \[x=\log_{10}\left( 41/9 \right) \]

OpenStudy (anonymous):

10 raised to the x-2

OpenStudy (anonymous):

Thanks!